If x, y ∈ Q then there are integers a, b, c such that x = a/c and y =b/c
Can this be proven using a proof or is there a solid counterexample?
.

Answer by richard1234(7193)

Statement is true.

x and y are rational by definition, so let c equal the least common denominator of x and y. Then you can find a and b such that x = a/c and y = b/c.

Formally, let and where p, q, r, s are integers, and , (i.e. fractions in simplest form). Then , and is an example set of values.